Operations Report – March 14th

Crew 206 Operations Report 14 – MAR – 2019

SOL: 18

Name of person filing report: Gaspard Thieulin

Non-nominal systems: Generator

Notes on non-nominal systems: Radiator hoses and
alternator need replacement on generator. A contractor has
been arranged to make these repairs within a few days.

Generator (hours run): 11,5 hours

Solar— SOC % – Turned off at 100%, Still off at 93% (1930)

Diesel Reading – 60%
Propane Reading – 72%

Ethanol Free Gasoline – 2 gallons

Water (Auxillary tank) – Not in use

Water (Static tank) – 315 gallons

Water differential (static tank) – 16 gallons

Gallons transferred: Not Applicable

Water in GreenHab – 204 gallons

Static to Loft Pump used – Yes

Water Meter: 0141772,2 gallons

Toilet tank emptied: Yes

Deimos rover used: No, still not functional

Hours: N/A

Beginning charge: N/A

Ending charge: N/A

Currently charging: N/A

Sojourner rover used: ASSIGNED TO DIRECTOR

Hours: Not Applicable

Beginning charge: Not Applicable

Ending charge: Not Applicable

Currently charging: Not Applicable

Opportunity rover used: Breaking system being repared.

Hours: 53,1 hours

Beginning charge: 100%

Ending charge: 100%

Currently charging: No

Curiosity rover used: Used

Hours: 81,3 hours

Beginning charge: 100%

Ending charge: 84%

Currently charging: Yes

Spirit rover used: Used

Hours: 82,7 hours

Beginning charge: 100%

Ending charge: 73%

Currently charging: Yes

Notes on rovers: Opportunity’s brake fluid is full again, we will check
its status on tomorrow’s EVA. Curiosity has been tested all day, no
particular issue has bee, discovered yet. We continue working on it.

ATV’s Used: N/A

Reason for use: Not used

Oil Added? N/A

ATV Fuel Used: N/A

# Hours the ATVs were Used today: None

Notes on ATVs: None

Hab Car used and why, where? Used by Staff.

Staff Crew Car used and why, where? Off Campus

General notes and comments: Nothing to report.

Summary of internet: Nothing to report.

Summary of suits and radios: Nothing to report.

Summary of Hab operations: Nothing to report.

Summary of Green Hab operations: Nothing to report.

Summary of Science Dome operations: Nothing to report.

Summary of RAM operations: Nothing to report.

Summary of any observatory issues: Nothing to report.

Summary of health and safety issues: See HSO Report.

Questions, concerns and requests to Mission Support: Nothing more to report